# Lower semicontinuity, Stoilow factorization and principal maps

### v2

@@ 1.2. Main results
Notice that equi-integrability of the negative part in (1.14) is needed even in the case of the determinant.
In this introduction we have stated the homogeneous versions of the lower semicontinuity estimates, but in the bulk of the text we will deal with general non autonomous $z$-dependent versions of the integrands. Before we embark on the proofs we present examples of principal quasiconvex functionals in section 2. In section 3 we describe the proofs. The proofs of Theorems 1.8 and 1.6 are of the same nature. First we use localization, we then deal with the gradient Young measures via Stoilow factorization, and finally we apply principal quasiconvexity. The localization for the case of asymptotic finite distortion is however rather delicate, as for example the 0-1 law from [3] is used. The proof of Theorem 1.5 is more subtle as we cannot rely on gradient Young measure theory. Instead we analyze the limit measure of the $\mathrm{L}^{1}$ bounded sequence $\mathbf{B}_{p}(\textup{D}u_{j})$ and localize differently at points where the limit measure is absolutely continuous with respect to Lebesgue measure or singular. This blow up technique adjusts very well to principal quasiconvexity, though to treat the singular part we need to assume critical integrability for the limit map (not for the sequence though).
@@ 2. Examples of Principal quasiconvex functionals
In this section we give examples of functionals satisfying the condition of Definition 1.2. A first simple class of examples is furnished by the following result:
@@
Assume $\Phi\colon\mathbb{R}^{2\times 2}\times\mathbb{R}\to\mathbb{R}\cup\{+\infty\}$ is convex and that the partial function $t\mapsto\Phi(A,t)$ is non-increasing for each fixed $A\in\mathbb{R}^{2\times 2}$. Then the functional $\mathbf{F}(A)\equiv\Phi(A,\det A)$, $A\in\mathbb{R}^{2\times 2}_{+}$, is $\mathrm{W}^{1,2}$-principal quasiconvex.
The proof follows easily if we combine Jensen’s inequality for convex functions with the area formula for principal maps of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ (see [5, Th. 2.10.1]). The details are left to the interested reader and we only remark here that a prime example of a polyconvex functional that is also $\mathrm{W}^{1,2}$-principal quasiconvex is the distortion functional $\mathbf{K}(A)\equiv K_{A}\equiv|A|^{2}/\det A$ for $\det A>0$.
@@
We now turn to the example from which all our research emanates, namely the well-known Burkholder functional $\mathbf{B}_{p}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}$ that we consider for exponents $p\geq 2$. Recall that in complex notation it is given as
@@
for a matrix $A\in\mathbb{R}^{2\times 2}$ that has conformal-anticonformal coordinates $(a_{+},a_{-})\in\mathbb{C}^{2}$. It is arguably the prime example of a rank-one convex functional candidate to be quasiconvex; in general, rank-one convexity is not sufficient for quasiconvexity [45]. On one hand, the Burkholder functional is structurally simple, since it is isotropic (that is, invariant under pre- and post-composition with rotations) and $p$-homogeneous. On the other, it is tightly linked to the optimal higher integrability properties of quasiconformal mappings, see [2]. In the recent work, [4] it was shown that the Burkholder functional is quasiconvex when restricted to quasiconformal test functions with appropriate distortion. In other words, the local Burkholder functional $\mathbf{B}_{K}$ defined in the introduction is quasiconvex. As described in [4], when taking the limit $p$ going to $2$ this quasiconvexity has interesting consequences also for existence of minimizers in elasticity.
In [4] it was proven that both $W(A)=K_{A}-\log K_{A}+\log(\det A)$ and the local Burkholder functional are principal quasiconvex. Our next result pinpoints which are the properties of the Burkholder functional which lead to its principal quasiconvexity:
@@
Let $p>2$ and $\mathbf{F}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}$ be a $p$-homogeneous, isotropic and quasiconvex functional on $Q_{2}(K)$ with $K\equiv p/(p-2)$. The quasiconvexity condition required is more precisely that
@@
holds for all $A\in Q_{2}(K)$ and $\phi\in\mathrm{W}^{1,2}_{0}(\mathbb{D})$ with $A+\textup{D}\phi(z)\in Q_{2}(K)$ a.e.
Assume $\mathbf{F}(A)\leq 0$ for all $A\in Q_{2}(K)$ and define the local functional
@@
$\mathbf{F}_{K}(A)\equiv\left\{\begin{array}[]{ll}\mathbf{F}(A)&\mbox{ if }A\in Q_{2}(K)\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus Q_{2}(K).\end{array}\right.$

### v3

@@ 1.2. Main results
Notice that we have not imposed any condition on the value of $\bf E$} at 0, so any value for ${\bf E}(0)$ that is compatible with lower semicontinuity of $\bf E$} is permitted in Theorem 1.9.
In this introduction we have stated the homogeneous versions of the lower semicontinuity estimates, but in the bulk of the text we will deal with general non-autonomous $z$-dependent versions of the integrands. Although we have stated the results for planar maps, the reader interested in applications could think of higher dimensional plane strains, which often can be modeled by two-dimensional deformations [20, 73, 74]. For example this could be relevant for isochoric Mooney-Rivlin materials altered in a quasiconvex but non-polyconvex way; see [24, 79] and the references therein. We remark that our results in particular yield semicontinuity results for strain energies that decompose as a sum of a polyconvex energy density and a principally quasiconvex energy density. This is particularly relevant for a class of rank-one convex functionals with an additive isochoric-volumetric split, see [84, 85] and [5, Sect. 11]. We also remark that under natural coercivity conditions in terms of either the distortion or the second invariant one obtains existence of minimizers subject to, for instance, Dirichlet boundary conditions. See [5, Sect. 12].
Before we embark on the proofs, we discuss and present examples of principal quasiconvex functionals in section 2. We start by recalling our earlier work on the Burkholder functional and isotropic functionals that admit a volumetric isochoric split. Additionally, we show that for any functional depending only on distortion, principal quasiconvexity is equivalent to rank-one convexity and to polyconvexity. We then investigate the natural question of which additional conditions, on top of quasiconvexity, imply principal quasiconvexity. Here, Proposition 2.6 is the main technical result. The proof originates from [5], utilizing the theory of quadrature domains, but in this work we deal more efficiently with possible concentration effects. This part combines the approximation result [50, 28] together with some careful manipulations with Orlicz functions in order to transfer equiintegrability from $K_{f}$ to $\log(1+\frac{1}{J_{f}})$, in analogy with the $\mathrm{L}^{p}$ estimates of Koskela and Onninen. More precisely, if $\bf E$} is $\mathrm{W}^{1,\infty}$-quasiconvex, satisfies a natural $\mathrm{L}^{1}$ type growth condition, see (2.8) in Section 2, and if for each $A\in\mathbb{R}^{2\times 2}_{+}$ the function
(1.22) $\mathbb{C}\setminus\{0\}\ni w\mapsto\mathbf{E}(wA)$
is superharmonic, then $\bf E$} is also $\mathrm{W}^{1,1}$-principal quasiconvex. For isotropic functionals the condition (1.22) simplifies further to an ordinary differential inequality, see Lemma 2.7. We end the examples section with a brief discussion of the so-called complex Burkholder functionals introduced in [12]. It is interesting to note that these functionals have a different symmetry group, but still the arguments persist in this wider context.
In section 3 we describe the proofs for the semicontinuity and existence results. The proofs of Theorems 1.6 and 1.9 follow a well-known strategy. First we use Young measures for the localization, we then deal with the homogeneous gradient Young measures via Stoilow factorization, and finally we apply principal quasiconvexity. The localization for the case of asymptotic finite distortion is however rather delicate, and involves a number of technical steps and results, where for instance the 0-1 law from [4] is used. The proof of Theorem 1.5 is more subtle as we cannot rely on gradient Young measure theory. Instead we analyze the limit measure of the $\mathrm{L}^{1}$-bounded sequence $\mathbf{B}_{p}(\textup{D}u_{j})$ and localize differently at points where the limit measure is absolutely continuous with respect to Lebesgue measure, and at points where it is singular. This blow up technique adjusts very well to principal quasiconvexity, though to treat the singular part we need to assume critical integrability for the limit map (not for the sequence though). Finally, we gather all results and prove in this section the results of existence of minimizers. The reader not interested in proofs can directly check our local conditions and apply the existence theorem.
@@ 2. Examples of principal quasiconvex functionals
In this section we discuss and give examples of functionals satisfying the principal quasiconvexity condition of Definition 1.2. Some of the examples are better understood using complex notation for matrices $A\in\mathbb{R}^{2\times 2}$. Recall that this amounts to identifying $A$ with $(a_{+},a_{-})\in\mathbb{C}^{2}$, where, in terms of the usual identification of $\omega=(\omega_{1},\omega_{2})\in\mathbb{R}^{2}$ with $\omega=\omega_{1}+\mathrm{i}\omega_{2}\in\mathbb{C}$,
(2.1) $A\omega=a_{+}\omega+a_{-}\overline{\,\omega\,}.$
Accordingly we get with $A^{+}\equiv(A+\mathrm{cof}A)/2$, $A^{-}\equiv(A-\mathrm{cof}A)/2$ that $A^{+}=(a_{+},0)$, $A^{-}=(0,a_{-})$ and this is the reason for calling $(a_{+},a_{-})$ the conformal-anticonformal coordinates of $A$. In this connection, we also record that $(\rho\mathrm{e}^{\mathrm{i}\theta},0)=\rho R_{\theta}$, where $R_{\theta}\in\mathrm{SO}(2)$ represents the (anti-clockwise) $\theta$-angle rotation about the origin. Similarly, the complex conjugate mapping, $\omega\mapsto\overline{\omega}$, is identified with the reflection $\mathrm{diag}(1,-1)=(0,1)$.
For later use, we also recall that for $A\in\mathbb{R}^{2\times 2}$ we have the identities: $|A|=|a_{+}|+|a_{-}|$, $J_{A}=|a_{+}|^{2}-|a_{-}|^{2}$. When $A\in\mathbb{R}^{2\times 2}_{+}$ we have $K_{A}=(|a_{+}|+|a_{-}|)/(|a_{+}|-|a_{-}|)=(1+|\mu_{A}|)/(1-|\mu_{A}|)$, where the ratio $\mu_{A}\equiv a_{-}/a_{+}$ is called the complex dilation of $A$.
A functional ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is isotropic provided ${\bf E}(QAR)={\bf E}(A)$ holds for all $A\in\mathbb{R}^{2\times 2}_{+}$ and $Q$, $R\in\mathrm{SO}(2)$. There are a number of equivalent ways to characterize isotropy and here we emphasize that both $A\mapsto J_{A}$ and $A\mapsto K_{A}$ are isotropic. These two functionals can be considered the building blocks for all isotropic functionals in the sense that one can show $\bf E$} to be isotropic precisely when ${\bf E}(A)=E(K_{A},J_{A})$ holds for all $A\in\mathbb{R}^{2\times 2}_{+}$ for some extended real-valued function $E\colon(0,\infty)^{2}\to\mathbb{R}\cup\{+\infty\}$. Below we shall be considering subclasses of the isotropic functionals that are naturally expressed in such terms. Finally, it will be convenient on a few occasions to use the short-hand $zA$ for the matrix $\rho R_{\theta}A=(z,0)(a_{+},a_{-})=(za_{+},za_{-})$, where $z=\rho\mathrm{e}^{\mathrm{i}\theta}$. For later reference we also record the general multiplication rule for matrices $A=(a_{+},a_{-})$ and $B=(b_{+},b_{-})$ in complex notation
(2.2) $AB=(a_{+},a_{-})(b_{+},b_{-})=\bigl{(}a_{+}b_{+}+a_{-}\overline{b_{-}},a_{+}b_{-}+a_{-}\overline{b_{+}}\bigr{)}.$
A first simple class of examples is furnished by the following result:
@@
Assume $P\colon\mathbb{R}^{2\times 2}\times\mathbb{R}\to\mathbb{R}\cup\{+\infty\}$ is an extended real-valued, lower semicontinuous and convex function for which the partial function $t\mapsto P(A,t)$ is non-increasing on $0,|A|^{2}]$ for each fixed $A\in\mathbb{R}^{2\times 2}$. Then the functional $\mathbf{P}(A)\equiv P(A,J_{A})$, $A\in\mathbb{R}^{2\times 2}_{+}$, is $\mathrm{W}^{1,1}$-principal quasiconvex.
Proof.
Let $f\colon\mathbb{C}\to\mathbb{C}$ be an orientation-preserving homeomorphism of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\mathbb{C})$, which is conformal on $\mathbb{C}\setminus\overline{\mathbb{D}}$ with the Laurent expansion
(2.3) $f(z)=b_{0}z+\frac{b_{1}}{z}+\sum_{n=2}^{\infty}\frac{b_{n}}{z^{n}}\,\mbox{ for }\,|z|>1.$
Note that a principal map in the sense of our Definition 1.1 in particular will satisfy these conditions. According to the area formula (see [8, Th. 2.10.1 & Cor. 2.10.2]) we have, in terms of the notation introduced after Definition 1.1, that
(2.4) $\fint_{\mathbb{D}}\!J_{f}\,\mathrm{d}m(z)\leq\det A_{f}-\sum_{n=2}^{\infty}n|b_{n}|^{2},$
and therefore in particular, $0\leq\fint_{\mathbb{D}}\!J_{f}\,\mathrm{d}m(z)\leq\det A_{f}$. We employ this inequality in conjunction with Jensen’s inequality whereby
$\fint_{\mathbb{D}}\!\mathbf{P}(\textup{D}f)\,\mathrm{d}m(z)\geq P\left(A_{f},\fint_{\mathbb{D}}\!J_{f}\,\mathrm{d}m(z)\right).$
Here we have $0\leq\fint_{\mathbb{D}}\!J_{u}\,\mathrm{d}m(z)\leq\det A_{f}\leq|A_{f}|^{2}$, where the last inequality is Hadamard’s. We may now conclude with the desired inequality if we use that $t\mapsto P(A_{f},t)$ is non-increasing on $0,|A_{f}|^{2}]$. ∎
We record that the above argument also illustrates why $A\mapsto J_{A}$ is not $\mathrm{W}^{1,\infty}$-principal quasiconvex. Indeed the area inequality (2.4) immediately implies that $\fint_{\mathbb{D}}\!J_{f}\,\mathrm{d}m(z)<\det A_{f}$ whenever $f$ is a $\mathrm{W}^{1,\infty}_{\operatorname{loc}}$ principal map where at least one of the coefficients $b_{n}\neq 0$ for $n\geq 2$ in its Laurent expansion (2.3). Prime examples of polyconvex functionals that are also $\mathrm{W}^{1,1}$-principal quasiconvex are, minus the determinant, $J_{A}$, and the distortion functional $\mathbf{K}(A)\equiv K_{A}$. It is well-known from [19] that $A\mapsto-J_{A}$ is $\mathrm{W}^{1,p}$-quasiconvex if and only if $p\geq 2$, so the fact that we in the definition of $\mathrm{W}^{1,1}$-principal quasiconvexity only test with special $\mathrm{W}^{1,1}$ maps is essential. See also [69] for further discussion of the $\mathrm{W}^{1,p}$-quasiconvexity condition that involves the distributional Jacobian.
The second invariant $\mathbf{I_{2}}(A)=K_{A}+1/K_{A}$ is also simultaneously polyconvex and $\mathrm{W}^{1,1}$-principal quasiconvex. This is a consequence of the following result concerning so-called isochoric, or conformally invariant, free energy functionals, meaning those isotropic functionals that are also homogeneous of degree 0. The canonical example of an isochoric functional is the distortion $A\mapsto K_{A}$ and it can be shown that ${\bf H}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is isochoric precisely when ${\bf H}(A)=H(K_{A})$ for all $A\in\mathbb{R}^{2\times 2}_{+}$, where $H\colon[1,\infty)\to\mathbb{R}\cup\{+\infty\}$. It is sometimes convenient to use the equivalent characterization in terms of the second invariant. Namely, $\bf H$} is isochoric precisely when ${\bf H}(A)=\tilde{H}({\bf I}_{A})$, where $\tilde{H}\colon[2,\infty)\to\mathbb{R}\cup\{+\infty\}$. For instance, see [84] and the references therein. This class of energy functionals and the associated variational problems have been discussed by a number of authors, including [8, Chap. 21], [9], [64] and [84, 85, 86]. The following result is stated in terms of the distortion function, but an entirely analoguous statement can be given in terms of the second invariant. We leave the details of this to the interested reader.
Proposition 2.2.
Let $H\colon[1,\infty)\to\mathbb{R}\cup\{+\infty\}$ be a lower semicontinuous and extended real-valued function. Define ${\bf H}(A)\equiv H(K_{A})$, $A\in\mathbb{R}^{2\times 2}_{+}$. Then the following are equivalent:
  • (i)
    $\bf H$} is rank-one convex;
  • (ii)
    $\bf H$} is $\mathrm{W}^{1,1}$-principal quasiconvex;
  • (iii)
    $\bf H$} is polyconvex;
  • (iv)
    $H$ is convex and non-decreasing.
Proof.
The pairwise equivalences of (i), (iii) and (iv) follow from [84]. To see the equivalence to (ii) we first observe that an argument similar to the one used in the proof of Proposition 2.1 yields (ii) from (iv). We therefore conclude the proof if we can show that (i) follows from (ii). Assume that (ii) holds and fix matrices $A_{0}$, $A_{1}\in\mathbb{R}^{2\times 2}_{+}$ that are rank-one connected. Put for a weight $\lambda\in(0,1)$, $A_{\lambda}=(1-\lambda)A_{0}+\lambda A_{1}$. We can without loss in generality assume that ${\bf H}(A_{0})$, ${\bf H}(A_{1})<+\infty$. Note that $A_{\lambda}\in\mathbb{R}^{2\times 2}_{+}$ and $1\leq K_{A_{i}}<+\infty$, hence by rank-one convexity of the distortion, $K_{A_{\lambda}}\leq(1-\lambda)K_{A_{0}}+\lambda K_{A_{1}}\leq K\equiv\max\{K_{A_{0}},K_{A_{1}}\}<+\infty$. Write $A_{1}-A_{0}=a\otimes n$ with $|n|=1$, take $\theta\in\mathrm{SO}(2)$ with $\theta e_{1}=n$, denote $Q=(0,1)^{2}$ and $R=\theta\bigl{(}Q\bigr{)}$. Define, in terms of real notation,
$v(z)\equiv\left\{\begin{array}[]{ll}(A_{0}-A_{\lambda})z+(1-\lambda)a&\mbox{ in }R_{0},\\ (A_{1}-A_{\lambda})z&\mbox{ in }R_{1},\end{array}\right$.
where $R_{i}=\theta\bigl{(}Q_{i}\bigr{)}$, $Q_{0}=(0,1-\lambda)\times(0,1)$, $Q_{1}=(1-\lambda,1)\times(0,1)$. Observe that hereby $v\colon R\to\mathbb{R}^{2}$ is a Lipschitz map that is constant in the direction perpendicular to $n$ and $v\bigl{(}\theta(te_{2})\bigr{)}=v\bigl{(}n+\theta(te_{2})\bigr{)}$ for $t\in(0,1)$. If therefore we extend $v$ to $\mathbb{R}^{2}$ by $R$-periodicity, then we obtain a Lipschitz and $R$-periodic map that we again denote $v\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$. By inspection
[…]
hence $\int_{R}\!\textup{D}v\,\mathrm{d}m(z)=0$. Put $u_{j}(z)\equiv A_{\lambda}z+\frac{1}{j}v(jz)$, $z\in\mathbb{D}$, and record that $u_{j}$ is Lipschitz with
[…]
where the disjoint sets $R_{0,j}$, $R_{1,j}\subset\mathbb{D}$ satisfy $\mathscr{L}^{2}(R_{0,j})=(1-\lambda)\pi+\mathrm{o}(1)$, $\mathscr{L}^{2}(R_{1,j})=\lambda\pi+\mathrm{o}(1)$. In particular, $K_{u_{j}}\leq K$, so $u_{j}$ is $K$-quasiregular and we may Stoilow factorize it: $u_{j}=h_{j}\circ f_{j}$ on $\mathbb{D}$, where $f_{j}\colon\mathbb{C}\to\mathbb{C}$ is a $K$-quasiconformal normalized principal map with $K_{f_{j}}=K_{u_{j}}$ and $h_{j}\colon f_{j}(\mathbb{D})\to\mathbb{C}$ is holomorphic. It follows from the Riemann-Lebesgue lemma that $u_{j}\overset{\ast}{\rightharpoonup}A_{\lambda}$ in $\mathrm{W}^{1,\infty}(\mathbb{D})$ and so by uniqueness of the Stoilow factorization we find that $f_{j}\to f$ uniformly on $\mathbb{C}$ and weakly in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and $h_{j}\to h$ locally uniformly on $f(\mathbb{D})$. Here we have in terms of $A_{\lambda}=(a_{\lambda,+},a_{\lambda,-})$ and in complex notation,
$f(z)=\left\{\begin{array}[]{ll}z+\frac{a_{\lambda,-}}{a_{\lambda,+}}\bar{z}&\mbox{ if }|z|\leq 1,\\ z+\frac{a_{\lambda,-}}{a_{\lambda,+}}\frac{1}{z}&\mbox{ if }|z|>1,\\ \end{array}\right.\quad\mbox{ and }\quad h(z)=a_{\lambda,+}z.$
Because
$K_{f_{j}}=K_{u_{j}}=\left\{\begin{array}[]{l}K_{A_{0}}\mbox{ on }R_{0,j},\\ K_{A_{1}}\mbox{ on }R_{1,j},\end{array}\right$.
we get
$\fint_{\mathbb{D}}\!{\bf H}(\textup{D}f_{j})\,\mathrm{d}m(z)=(1-\lambda){\bf H}(A_{0})+\lambda{\bf H}(A_{1})+\mathrm{o}(1)$
and here the left-hand side is estimated using $\mathrm{W}^{1,1}$-principal quasiconvexity,
${\bf H}\left(\fint_{\mathbb{D}}\!\textup{D}f_{j}\,\mathrm{d}m(z)\right)\leq\fint_{\mathbb{D}}\!{\bf H}(\textup{D}f_{j})\,\mathrm{d}m(z).$
Because $\fint_{\mathbb{D}}\!\textup{D}f_{j}\,\mathrm{d}m(z)\to A_{f}=A_{\lambda}/a_{\lambda,+}$ and since $K_{A_{\lambda}/a_{\lambda,+}}=K_{A_{\lambda}}$ we may use the lower semicontinuity of $H$ to conclude. ∎
Next we turn to the class of free energy functionals that are isotropic and admit an additive isochoric-volumetric split, meaning they must have the general form
${\bf E}(A)=H(K_{A})+G(J_{A}),\quad A\in\mathbb{R}^{2\times 2}_{+},$
where $H$ and $G$ are extended real-valued functions. Such free energy functionals have been considered by many authors and are often used to model slightly compressible materials, see for instance [38, 30, 47, 73]. Whereas the case of rank-one convexity for the individual terms $H(K_{A})$ and $G(J_{A})$ is characterized by Propositions 2.2 and 2.1, respectively, the rank-one convexity of ${\bf E}(A)$ is more involved and does not necessarily mean that both $H(K_{A})$ and $G(J_{A})$ are rank-one convex. The precise conditions have been worked out in the interesting paper [84, Th. 2.6] and following [85] we shall here focus on the special case when the isochoric part is rank-one convex. In that situation it was shown in [85] (see also [5, Sect. 11]) that $\bf E$} is rank-one convex if and only if there exists an isotropic polyconvex functional $\bf P$} and a non-negative constant $c\geq 0$ such that ${\bf E}={\bf P}+c{\bf W}$, where
(2.5) ${\bf W}(A)\equiv K_{A}-\log K_{A}+\log J_{A}\,\mbox{ for }A\in\mathbb{R}^{2\times 2}_{+},$
as observed in [85], is an isotropic rank-one convex, non-polyconvex functional. Indeed it cannot be polyconvex because ${\bf W}(tA)\to$-$\infty$ as $t\to 0$. If we follow the convention (1.12) of this paper and extend $\bf W$} to full matrix space $\mathbb{R}^{2\times 2}$ and at the same time insist that the extension be lower semicontinuous, then we must define
${\bf W}(A)=\left\{\begin{array}[]{ll}K_{A}-\log K_{A}+\log J_{A}&\mbox{ if }A\in\mathbb{R}^{2\times 2}_{+},\\ -\infty&\mbox{ if }A=0,\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus\bigl{(}\mathbb{R}^{2\times 2}_{+}\cup\{0\}\bigr{)}.\end{array}\right$.
We are therefore forced to consider extended real-valued integrands that assume both $\pm\infty$. This example, together with the local Burkholder functional (1.14), are our reasons for adopting an ad hoc approach to the value at 0 when extending functionals from $\mathbb{R}^{2\times 2}_{+}$ to $\mathbb{R}^{2\times 2}$. We remark that this extension of $\bf W$} remains isotropic and rank-one convex (with the natural definitions, see for instance [5]). The following result is proved in [5, Cor. 1.7, Th. 1.8 & Th. 1.9]. We emphasize that it in particular yields a positive solution to Morrey’s problem within the considered class of functionals.
Theorem 2.3.
Let ${\bf E}(A)=H(K_{A})+G(J_{A})$, $A\in\mathbb{R}^{2\times 2}_{+}$, where $H\colon[1,\infty)\to\mathbb{R}\cup\{+\infty\}$ and $G\colon(0,\infty)\to\mathbb{R}\cup\{+\infty\}$ are lower semicontinuous extended real-valued functions. Assume $H$ is convex. Then $\bf E$} is rank-one convex if and only if $\bf E$} is $\mathrm{W}^{1,2}$-quasiconvex.
Furthermore, the functional $\bf W$} from (2.5) is $\mathrm{W}^{1,1}$-principal quasiconvex.
@@
We return to the Burkholder functional $\mathbf{B}_{p}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}$ that was defined for exponents $p\geq 2$ at (1.13) in the introduction. In complex notation it is given by
@@
for a matrix $A\in\mathbb{R}^{2\times 2}$ that has conformal-anticonformal coordinates $(a_{+},a_{-})\in\mathbb{C}^{2}$. It is isotropic, $p$-homogeneous and rank-one convex, but it is an open question whether it is also $\mathrm{W}^{1,\infty}$-quasiconvex. Its quasiconvexity would among other things provide an alternative approach to the optimal higher integrability properties of quasiconformal maps [3], see [52] where the reader can also find interesting extensions to higher dimensions $n>2$. In the recent work, [5], building on [10, 11], it was shown that the Burkholder functional is quasiconvex when restricted to quasiconformal test functions with appropriate distortion. The precise result is conveniently stated in terms of the local Burkholder functional.
@@
Let $p>2$, $p^{\prime}=p/(p-1)$ and K=p/(p2)K=p/(p-2). The local Burkholder functional $\mathbf{B}_{K}^{\operatorname{loc}}$ defined at (1.14) is simultaneously $\mathrm{W}^{1,1}$-principal quasiconvex, $\mathrm{W}^{1,p^{\prime}}$-quasiconvex and closed $\mathrm{W}^{1,q}$-quasiconvex for each $q>p^{\prime}$.
We refer to [5, Th. 1.4 & Th. 1.6] for the proofs and note that the statement is essentially sharp. Indeed, if we test $\mathbf{B}_{K}^{\operatorname{loc}}$ with the staircase laminate introduced in [35, Example 5.4] we see that it is not closed $\mathrm{W}^{1,q}$-quasiconvex for $q<2K/(K+1)$. Moreover, if we test $\mathbf{B}_{K}^{\operatorname{loc}}$ with the map obtained in [7, Theorem 3.2] we see that it is not even $\mathrm{W}^{1,q}$-quasiconvex for $q<2K/(K+1)$. That $\mathbf{B}_{K}^{\operatorname{loc}}$ can be $\mathrm{W}^{1,1}$-principal quasiconvex is because we only require Jensen’s inequality for homeomorphisms in the definition of principal quasiconvexity and a quasiconformal homeomorphism of Sobolev class $\mathrm{W}^{1,1}_{\operatorname{loc}}$ is automatically in $\mathrm{W}^{1,2}_{\operatorname{loc}}$. [^2] This is classical. A proof follows by applying, for example [8, Corollary 3.3.6], to get that the Jacobian is in $\mathrm{L}^{1}$ and deduce the square integrability of the full derivative by the distortion inequality.
We have already noted that $\mathbf{B}_{K}^{\operatorname{loc}}$ is non-positive on its effective domain which is the $K$-quasiconformal well $Q_{2}(K)$. Indeed, it is 0 on its boundary and strictly negative in its interior. If we combine Theorem 2.4 with Jensen’s inequality for convex functions we deduce the following corollary.
Corollary 2.5.
Let $p>2$, $p^{\prime}=p/(p-1)$ and K=p/(p2)K=p/(p-2). Assume $\theta\colon(-\infty,0)\to\mathbb{R}\cup\{+\infty\}$ is a lower semicontinuous convex and non-decreasing function satisfying $\theta(t)\to+\infty$ as $t\to 0$-. Then the functional $\mathbf{B}_{p}^{\theta}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ defined by
@@
is $\mathrm{W}^{1,1}$-principal quasiconvex, $\mathrm{W}^{1,p^{\prime}}$-quasiconvex and closed $\mathrm{W}^{1,q}$-quasiconvex for each $q>p^{\prime}$.
We leave the elementary verification to the interested reader. We shall exhibit one particular example from this and start by recalling that $\mathbf{B}_{p}$ is isotropic. If therefore we restrict $\mathbf{B}_{p}(A)$ to $A\in\mathbb{R}^{2\times 2}_{+}$ it can be expressed in terms of the distortion and the Jacobian of $A$, where, if we assume $p>2$ and keep writing K=p/(p2)K=p/(p-2), it becomes
$\mathbf{B}_{p}(A)=\frac{p-2}{2}\bigl{(}K_{A}-K\bigr{)}K_{A}^{\frac{p-2}{2}}J_{A}^{\frac{p}{2}},\quad A\in\mathbb{R}^{2\times 2}_{+}.$
It clearly does not admit an additive isochoric-volumetric split. However, it does admit a multiplicative isochoric-volumetric split as
$\mathbf{B}_{p}(A)=\Phi(K_{A})\Psi(J_{A}),$
where $\Phi(s)=\frac{p-2}{2}(s-K)s^{\frac{p-2}{2}}$, $s\geq 1$, and $\Psi(t)=t^{\frac{p}{2}}$, $t>0$. It is interesting to note that both $\Phi$ and $\Psi$ are increasing and convex on their respective domains. We exploit this as follows and note first that $\Phi(s)<0$ for $s\in[1,K)$ corresponding to $\mathbf{B}_{p}(A)<0$ on the interior of the $K$-quasiconformal well. We may therefore define the functional $\mathbf{B}^{\theta}_{p}$ corresponding to $\theta(t)=-\log(-t)$, $t<0$, whereby the functional
@@
(2.7) $\mathbf{L}_{p}(A)\equiv\mathbf{B}_{p}^{\theta}(A)=\left\{\begin{array}[]{ll}-\log\bigl{(}-\Phi(K_{A})\bigr{)}-\log\bigl{(}\Psi(J_{A})\bigr{)}&\mbox{ if }A\in\mathbb{R}^{2\times 2}_{+}\mbox{ and }K_{A}<K,\\ +\infty&\mbox{ otherwise,}\end{array}\right$.
results and it clearly has the additive isochoric-volumetric split form. By virtue of Propositions 2.1 and 2.2 both the isochoric and volumetric parts are polyconvex, so $\mathbf{L}_{p}$ is in particular polyconvex too.
The following proposition gives, modulo a one-sided growth condition, a local condition for a quasiconvex function to be automatically principal quasiconvex. The interested reader will see that the statement in particular applies to the functional $\bf W$} defined at (2.5).
Proposition 2.6.
Let ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}$ be $\mathrm{W}^{1,\infty}$-quasiconvex. Suppose that
(2.8) ${\bf E}(A)\leq c\left(1+K_{A}+\log\left(1+\frac{1}{J_{A}}\right)+J_{A}^{\alpha}\right)$
and that the function
(2.9) $\mathbb{C}\setminus\{0\}\ni w\mapsto{\bf E}(wA)\,\mbox{ is superharmonic}$